OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later.

OpenStudy (anonymous):

substitute x=6 in the given eq i.e. y = 65 - 2x = 65-2*6=65-12=53 i.e. y=53 Hence the temperature 6 hours later will be 53 degree Celsius.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

hey can you help me with this question please ? Find the direct variation equation which passes through (0, 0) and (4, 1)

OpenStudy (anonymous):

Since the points are (0, 0) and (4, 1) ,So, x1=0, x2=4, y1=0, y2=1 Therefore eq is given as \[\frac{x-x_{1}}{x_{1}-x_{2}} = \frac{y-y_{1}}{y_{1}-y_{2}}\] i.e. \[\frac{x-0}{0-4} = \frac{y-0}{0-1} \rightarrow \frac{x}{-4} = \frac{y}{-1} \] i.e. \[\frac{x}{-4} = \frac{y}{-1} \rightarrow y = \frac{1}{4} x \rightarrow y \alpha x\] Which is the required eq of direct variation.

OpenStudy (anonymous):

\[y=\frac{1}{4} x \] is the required ans